Calculate the pH of a buffer solution prepared by dissolving 19.50 g of sodium cyanate NaCNO (65.01g/mol) in 1.00 M cyanic acid (HCNO) to a final volume of 1.00 liter of solution.    Ka(HCNO) = 3.50 x 10-4
Answer #1

mass(NaCNO)= 19.50 g

use:

number of mol of NaCNO,

n = mass of NaCNO/molar mass of NaCNO

=(19.5 g)/(65.01 g/mol)

= 0.3 mol

volume , V = 1 L

use:

Molarity,

M = number of mol / volume in L

= 0.3/1

= 0.3 M

Ka = 3.5*10^-4

pKa = - log (Ka)

= - log(3.5*10^-4)

= 3.456

use:

pH = pKa + log {[conjugate base]/[acid]}

= 3.456+ log {0.3/1}

= 2.933

Answer: 2.93